OpenStudy (anonymous):

R1= 3 ohms R2= 6 ohms R3= 1 ohms V= 9 volts A. What is the total (net) resistance? B. What is the total current (in amps)? C. What is the current (in amps) across each resistor?

OpenStudy (anonymous):

are resistors in series or in parallel ?

OpenStudy (anonymous):

series. @ASAAD123

OpenStudy (anonymous):

A) R(net)=R1+R2+R3 B)&C) I=(9/R(net)) A